# On the Empirical Power of Goodness-of-Fit Tests in Watermark Detection

### v1

@@ 2 Preliminaries
Li et al. [31] extend it to a more practical setting where the generated text may be partially edited. They model human edits as a mixture effect so that $Y_{t}$ is from a mixture of $\mu_{0}$ and $\mu_{1,\bm{P}_{t}}$ under $H_{1}$.
@@ 3 GoF Tests for Watermark Detection
Tr-GoF test,
Phi, [31]
$S_{n}^{+}(s)=\sup\limits_{r\in[{\textnormal{p}}^{+},1)}K_{s}^{+}(F_{n}(r),F_{0}(r))$
Kuiper’s test,
Kui, [29]
$V_{n}=\max\limits_{1\leq i\leq n}\left(\frac{i}{n}-{\textnormal{p}}_{(i)}\right)+\max\limits_{1\leq i\leq n}\left({\textnormal{p}}_{(i)}-\frac{i-1}{n}\right)$
Kolmogorov-Smirnov test
Kol, [47]
$D_{n}=\max\limits_{1\leq i\leq n}\left\{\max\left({\textnormal{p}}_{(i)}-\frac{i-1}{n},\frac{i}{n}-{\textnormal{p}}_{(i)}\right)\right\}$
Anderson-Darling test,
And, [2]
$A^{2}=-n-\frac{1}{n}\sum_{i=1}^{n}(2i-1)\log\big[$
$\qquad\qquad(1-{\textnormal{p}}_{(n+1-i)}){\textnormal{p}}_{(i)}\big]$
Cramér-von Mises test,
Cra, [6]
$W^{2}=\frac{1}{12n}+\sum_{i=1}^{n}\left({\textnormal{p}}_{(i)}-\frac{2i-1}{2n}\right)^{2}$
Watson’s test,
Wat, [51]
$U^{2}=W^{2}-n\left(\bar{F}-\frac{1}{2}\right)^{2}$
Neyman’s smooth test,
Ney, [38]
$T_{k}=n\sum_{j=1}^{k}a_{j}^{2}$
Chi-squared test,
Chi, [40]
$\chi^{2}=\sum_{i=1}^{k}\frac{(O_{i}-E_{i})^{2}}{E_{i}}$
@@ 4.1 Experiment settings
Human edits can weaken watermark signals. To evaluate detection robustness, we apply two common types of edits: word deletion and synonym substitution. For each, we randomly select a fraction $r_{\text{edit}}\in\{0.1,0.2\}$ of watermarked tokens and either delete them or replace them with synonyms from WordNet [37].
@@ Acknowledgments
This work was supported in part by NIH grants R01AG071174, U01CA274576, R01EB036016, and P30AG073105, NSF grant DMS-2310679, a Meta Faculty Research Award, and Wharton AI for Business. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ A.3 SynthID Watermark
The SynthID watermark, proposed by Google [9], is based on a novel sampling strategy for categorical distributions called tournament sampling. Given a number of tournament rounds $k$, the pseudorandom seed is defined as $\zeta=(\bar{\bm{g}}_{i})_{i=1}^{k}$, where each $\bar{\bm{g}}_{i}=(g_{i,{w}})_{{w}\in{{\mathcal{W}}}}$ is a vector of i.i.d. samples drawn from U(0,1){U}(0,1). Unlike the previous two watermarking schemes, the decoder here is stochastic rather than deterministic. The next token is sampled from a modified NTP distribution, i.e., ${w}\sim{\mathcal{S}}^{\mathrm{syn}}(\bm{P},\zeta)$, where
${\mathcal{S}}^{\mathrm{syn}}(\bm{P},\zeta)={\mathcal{T}}_{\bar{\bm{g}}_{k}}\circ\cdots\circ{\mathcal{T}}_{\bar{\bm{g}}_{1}}(\bm{P}),$ (8)
where ${\mathcal{T}}_{\bar{\bm{g}}}$ is a vectorized operator defined as
$({\mathcal{T}}_{\bar{\bm{g}}}(\bm{P}))({w})=P_{{w}}\cdot\left(P_{{w}}+2\cdot\sum_{{w}^{\prime}:g_{{w}^{\prime}}<g_{{w}}}P_{{w}^{\prime}}\right).$ (9)
Dathathri et al. [9] prove that the output of ${\mathcal{T}}_{\bar{\bm{g}}}(\bm{P})$ corresponds to the distribution of the winner token in a one-versus-one competition, where two independent tokens sampled from $\bm{P}$ compete based on their assigned $\bar{\bm{g}}$ values. Repeating this process over $k$ rounds with i.i.d. pseudorandom vectors $\bar{\bm{g}}_{i}$ results in $\bm{P}_{\zeta}$ being the distribution of the final tournament winner after $k$ rounds of competition.
For detection, we recover the $g$-values $\zeta=(\bar{\bm{g}}_{i})_{i=1}^{k}$ and extract the coordinates corresponding to the selected token w{w}, i.e., $\{g_{i,{w}}\}_{i\in[k]}$. Under the tournament sampling scheme, these values tend to be larger under $H_{1}$ than under $H_{0}$, due to the winner-takes-all interpretation. We consider the sum-based detection rule proposed by Dathathri et al. [9] as our baseline, which computes a simple average over the selected $k$ $g$-values:

### v2

@@ 2 Preliminaries
Li et al. [31, 33] extend it to a more practical setting where the generated text may be partially edited. They model human edits as a mixture effect so that $Y_{t}$ is from a mixture of $\mu_{0}$ and $\mu_{1,\bm{P}_{t}}$ under $H_{1}$.
@@ 3 GoF Tests for Watermark Detection
Tr-GoF test, Phi, [31] $S_{n}^{+}(s)=\sup\limits_{r\in[{\textnormal{p}}^{+},1)}K_{s}^{+}(F_{n}(r),F_{0}(r))$ Kuiper’s test, Kui, [30] $V_{n}=\max\limits_{1\leq i\leq n}\left(\frac{i}{n}-{\textnormal{p}}_{(i)}\right)+\max\limits_{1\leq i\leq n}\left({\textnormal{p}}_{(i)}-\frac{i-1}{n}\right)$
Kolmogorov-Smirnov test Kol, [50] $D_{n}=\max\limits_{1\leq i\leq n}\left\{\max\left({\textnormal{p}}_{(i)}-\frac{i-1}{n},\frac{i}{n}-{\textnormal{p}}_{(i)}\right)\right\}$ Anderson-Darling test, And, [2] $A^{2}=-n-\frac{1}{n}\sum_{i=1}^{n}(2i-1)\log\big[$ $\qquad\qquad(1-{\textnormal{p}}_{(n+1-i)}){\textnormal{p}}_{(i)}\big]$
Cramér-von Mises test, Cra, [6] $W^{2}=\frac{1}{12n}+\sum_{i=1}^{n}\left({\textnormal{p}}_{(i)}-\frac{2i-1}{2n}\right)^{2}$ Watson’s test, Wat, [54] $U^{2}=W^{2}-n\left(\bar{F}-\frac{1}{2}\right)^{2}$
Neyman’s smooth test, Ney, [40] $T_{k}=n\sum_{j=1}^{k}a_{j}^{2}$ Chi-squared test, Chi, [42] $\chi^{2}=\sum_{i=1}^{k}\frac{(O_{i}-E_{i})^{2}}{E_{i}}$
@@ 4.1 Experiment settings
Human edits can weaken watermark signals [48, 19, 31]. To evaluate detection robustness, we apply two common types of edits: word deletion and synonym substitution. For each, we randomly select a fraction $r_{\text{edit}}\in\{0.1,0.2\}$ of watermarked tokens and either delete them or replace them with synonyms from WordNet [39].
@@ Acknowledgments
This work was supported in part by NIH grants R01AG071174, U01CA274576, R01EB036016, R01EB037101, U01AG066833, R01LM014731, and P30AG073105, NSF grant DMS-2310679, a Meta Faculty Research Award, and Wharton AI for Business. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ A.3 SynthID Watermark
The SynthID watermark, proposed by Google [9], is based on a novel sampling strategy for categorical distributions called tournament sampling. Given a number of tournament rounds $k$, the pseudorandom seed is defined as $\zeta=({\bm{g}}_{i})_{i=1}^{k}$, where each ${\bm{g}}_{i}=(g_{i,{w}})_{{w}\in{{\mathcal{W}}}}$ is a vector of i.i.d. samples drawn from U(0,1){U}(0,1). Unlike the previous two watermarking schemes, the decoder here is stochastic rather than deterministic. The next token is sampled from a modified NTP distribution, i.e., ${w}\sim{\mathcal{S}}^{\mathrm{syn}}(\bm{P},\zeta)$, where
${\mathcal{S}}^{\mathrm{syn}}(\bm{P},\zeta)={\mathcal{T}}_{{\bm{g}}_{k}}\circ\cdots\circ{\mathcal{T}}_{{\bm{g}}_{1}}(\bm{P}),$ (8)
where ${\mathcal{T}}_{{\bm{g}}}$ is a vectorized operator defined as
$({\mathcal{T}}_{{\bm{g}}}(\bm{P}))({w})=P_{{w}}\cdot\left(P_{{w}}+2\cdot\sum_{{w}^{\prime}:g_{{w}^{\prime}}<g_{{w}}}P_{{w}^{\prime}}\right).$ (9)
Dathathri et al. [9] prove that the output of ${\mathcal{T}}_{{\bm{g}}}(\bm{P})$ corresponds to the distribution of the winner token in a one-versus-one competition, where two independent tokens sampled from $\bm{P}$ compete based on their assigned $\bm{g}$} values. Repeating this process over $k$ rounds with i.i.d. pseudorandom vectors ${\bm{g}}_{i}$ results in $\bm{P}_{\zeta}$ being the distribution of the final tournament winner after $k$ rounds of competition.
For detection, we recover the $g$-values $\zeta=({\bm{g}}_{i})_{i=1}^{k}$ and extract the coordinates corresponding to the selected token w{w}, i.e., $\{g_{i,{w}}\}_{i\in[k]}$. Under the tournament sampling scheme, these values tend to be larger under $H_{1}$ than under $H_{0}$, due to the winner-takes-all interpretation. We consider the sum-based detection rule proposed by Dathathri et al. [9] as our baseline, which computes a simple average over the selected $k$ $g$-values: